# RaTrack: Moving Object Detection and Tracking with 4D Radar Point Cloud

### 10 Jan 2024 v3

@@ Abstract
Mobile autonomy relies on the precise perception of dynamic environments. Robustly tracking moving objects in 3D world thus plays a pivotal role for applications like trajectory prediction, obstacle avoidance, and path planning. While most current methods utilize LiDARs or cameras for Multiple Object Tracking (MOT), the capabilities of 4D imaging radars remain largely unexplored. Recognizing the challenges posed by radar noise and point sparsity in 4D radar data, we introduce RaTrack, an innovative solution tailored for radar-based tracking. Bypassing the typical reliance on specific object types and 3D bounding boxes, our method focuses on motion segmentation and clustering, enriched by a motion estimation module. Evaluated on the View-of-Delft dataset, RaTrack showcases superior tracking precision of moving objects, largely surpassing the performance of the state of the art.
@@ I Introduction
State-of-the-art techniques in moving object tracking or Multiple Object Tracking (MOT) primarily use on-vehicle LiDARs [10, 11, 12], cameras [13, 14, 15], or their fusion [16, 17, 18]. Surprisingly, the potential of 4D mmWave radars remains under-explored. As an emerging automotive sensor, 4D mmWave radar is gaining traction due to its improved imaging ability, resilience against challenging visual conditions (e.g., fog, rain, dust), ability to measure object velocities and cost-effectiveness. These merits make 4D radar an appealing and robust supplement, or even alternative, to automotive LiDARs.
However, integrating 4D radars into moving object tracking presents non-trivial challenges. The prevalent approaches [10, 11, 12, 19, 17] often follow the tracking-by-detection paradigm. Such a paradigm involves first detecting objects in each frame independently and then linking these detected object types and 3D bounding boxes across consecutive frames to form continuous object trajectories. Key to the tracking-by-detection success depends on the detection accuracy. This paradigm struggles when adapted to 4D radar data, due to the inherent radar noise and point sparsity, undermining accurate type classification and bounding box regression. Specifically, the non-negligible multi-path noises in radar data complicate the correct identification of objects while the sparsity of radar point clouds makes the object property (e.g. extension and orientation) regression even more difficult. As exhibited in [20], the mAP performance [21] of the 4D radar detection method is only 38.0, $\sim$60inferior to its LiDAR counterpart in the same scene. Such poor 3D detection performance compromises the reliability of 4D radar-based tracking in real-world scenarios.
To address this, we present RaTrack, a tailored solution for moving object tracking using 4D automotive radars. Our approach stems from a critical insight: for effective multi-object tracking, class-agnostic detection is often adequate, and the conventional reliance on 3D bounding boxes becomes redundant if distinct point clusters can be utilized. Driven by this understanding, we restructure the moving object detection challenge into simpler motion segmentation and clustering tasks. This restructuring allows us to sidestep the complex tasks of object type classification and bounding box regression typically encountered with 4D radars. To further enhance our method’s performance, we integrate a motion estimation module, which enriches the inherently sparse radar data with point-level scene flow. Building on this, our data association module is precisely adapted to our clustering method and is calibrated to weigh different features for optimal matching. Our solution is architected as an end-to-end trainable network, with its training modelled as a multi-task learning endeavor. This encompasses motion segmentation, scene flow estimation, and affinity matrix computation.
@@ III Problem Formulation
Notation. Given this context, we consider the problem of online moving object detection and tracking with 4D automotive radar. The input is an ordered 4D radar point cloud sequence $\mathcal{P}=\{\mathbf{P}^{t}\}_{t=1}^{T}$ comprised of $T$ frames captured by the same radar sensor mounted on a moving vehicle. A frame $\mathbf{P}^{t}=[\mathbf{p}^{t}_{1};...;\mathbf{p}^{t}_{i};...;\mathbf{p}^{t}_{N^{t}}]$ contains $N^{t}$ radar points. Each radar point $\mathbf{p}_{i}^{t}$ is characterised by its 3D position $\mathbf{x}_{i}^{t}\in\mathbb{R}^{3}$ in the metric space and auxiliary velocity features $\mathbf{v}_{i}^{t}=[v_{r,i}^{t},v_{c,i}^{t}]$, where $v_{r,i}^{t}$ and $v_{c,i}^{t}$ are the measured relative radial velocity (RRV) and its ego-motion compensated variant. Given each radar point cloud $\mathbf{P^{t}}$ from the stream, our objective is to detect multiple moving objects $\mathbf{D}^{t}=\{\mathbf{d}_{k}^{t}\}_{k=1}^{K^{t}}$ in a class-agnostic manner without the need to regress their 3D bounding boxes. These detected objects are then associated with objects tracked in the previous frame $\mathbf{O}^{t-1}=\{\mathbf{o}_{m}^{t-1}\}_{m=1}^{M^{t}}$. The result of this process is a set of updated objects $\mathbf{O}^{t}$ in track for the current frame.
@@ IV-A Overview
We introduce RaTrack, a generic learning-based framework bespoken for 4D radar-based moving object detection and tracking. As seen in Fig. 1, in our network pipeline, we first apply a backbone (c.f. Sec. IV-B) to encode intra- and inter-frame radar point cloud features. With the extracted features, our motion estimation module (c.f. Sec. IV-C) infers point-level scene flow as an explicit complement to augment the latent features of radar point clouds. Our advocated idea of class-agnostic detection without bounding boxes is introduced in the object detection module (c.f. Sec. IV-D), in which moving points are first identified and then used to detect moving objects via clustering. Finally, our data association module (c.f. Sec. IV-E) computes the affinity matrix with a learnable distance metric and then optimises the bipartite matching problem. The entire network is end-to-end trainable with a multi-task loss that incorporates three supervised subtasks: motion segmentation, scene flow estimation, and affinity matrix computation.
@@ IV-C Motion Estimation Module
Radar sensors’ inherent sparsity and noise result in latent features from radar point clouds that are deficient in informative geometric cues, complicating object detection and tracking. To address this, we introduce a motion estimation module that explicitly determines per-point motion vectors. Contrary to the conventional scene flow approach, which estimates per-point motion vectors in a forward direction (from frame $t$ to $t+1$), we opt for a backward estimation (from frame $t$ to $t-1$). This not only negates tracking latency but also ensures that the estimated flow vectors correspond to the points in the current frame $t$.
@@ V-A Evaluation Settings
Baselines. As there are no prior works for 4D radar-based moving object tracking, we select two state-of-the-art LiDAR-oriented 3D MOT methods, i.e., AB3DMOT [10] and CenterPoint [11] as our baselines. To ensure the comparison is fair, we keep their original settings and also train their models on the VoD training split. Specifically, we develop two augmented baselines AB3DMOT-PP and CenterPoint-PP by replacing the detector with PointPillars [51] for AB3DMOT and the backbone with that of PointPillars for CenterPoint. As our baselines detect and track 3D bounding boxes, we transform their object outputs into clusters of points so that their results can be compared with ours.

### 24 Feb 2024 v5

@@ Abstract
Mobile autonomy relies on the precise perception of dynamic environments. Robustly tracking moving objects in 3D world thus plays a pivotal role for applications like trajectory prediction, obstacle avoidance, and path planning. While most current methods utilize LiDARs or cameras for Multiple Object Tracking (MOT), the capabilities of 4D imaging radars remain largely unexplored. Recognizing the challenges posed by radar noise and point sparsity in 4D radar data, we introduce RaTrack, an innovative solution tailored for radar-based tracking. Bypassing the typical reliance on specific object types and 3D bounding boxes, our method focuses on motion segmentation and clustering, enriched by a motion estimation module. Evaluated on the View-of-Delft dataset, RaTrack showcases superior tracking precision of moving objects, largely surpassing the performance of the state of the art. We release our code and model at https://github.com/LJacksonPan/RaTrack.
@@ I Introduction
State-of-the-art techniques in moving object tracking or Multiple Object Tracking (MOT) primarily use on-vehicle LiDARs [10, 11, 12], cameras [13, 14, 15], or their fusion [16, 17, 18]. Surprisingly, the potential of 4D mmWave radars remains under-explored. As an emerging automotive sensor, 4D mmWave radar is gaining traction due to its improved imaging ability, resilience against challenging weather and illumination conditions (e.g., fog, dust, darkness, glare), ability to measure object velocities and cost-effectiveness. These merits make 4D radar an appealing and robust supplement, or even alternative, to automotive LiDARs.
However, integrating 4D radars into moving object tracking presents non-trivial challenges. The prevalent approaches [10, 11, 12, 19, 17] often follow the tracking-by-detection paradigm. Such a paradigm involves first detecting objects in each frame independently and then linking these detected object types and 3D bounding boxes across consecutive frames to form continuous object trajectories. Key to the tracking-by-detection success depends on the detection accuracy. This paradigm struggles when adapted to 4D radar data, due to the inherent radar noise and point sparsity, undermining accurate type classification and bounding box regression. Specifically, the non-negligible multi-path noises in radar data complicate the correct identification of objects while the sparsity of radar point clouds makes the object property (e.g. extension and orientation) regression even more difficult. As exhibited in [20], the mAP performance [21] of the 4D radar detection method is only 38.0, $\sim$40inferior to its LiDAR counterpart in the same scene. Such poor 3D detection performance compromises the reliability of 4D radar-based tracking in real-world scenarios.
To address this, we present RaTrack, a first-of-its-kind tailored solution for moving object tracking using 4D automotive radars. Our approach stems from a critical insight: for effective multi-object tracking, class-agnostic detection is often adequate, and the conventional reliance on 3D bounding boxes becomes redundant if distinct point clusters can be utilized. Driven by this understanding, we restructure the moving object detection challenge into simpler motion segmentation and clustering tasks. This restructuring allows us to sidestep the complex tasks of object type classification and bounding box regression typically encountered with 4D radars. To further enhance our method’s performance, we integrate a point-wise motion estimation module, which enriches the inherently sparse radar data with point-level scene flow. Building on this, our data association module is precisely adapted to our clustering method and is calibrated to weigh different features for optimal matching. Our solution is architected as an end-to-end trainable network, with its training modelled as a multi-task learning endeavour. This encompasses motion segmentation, scene flow estimation, and affinity matrix computation.
@@ III Problem Formulation
Notation. Given this context, we consider the problem of online moving object detection and tracking with 4D automotive radar. The input is an ordered 4D radar point cloud sequence $\mathcal{P}=\{\mathbf{P}^{t}\}_{t=1}^{T}$ comprised of $T$ frames captured by the same radar sensor mounted on a moving vehicle. A frame $\mathbf{P}^{t}=[\mathbf{p}^{t}_{1};...;\mathbf{p}^{t}_{i};...;\mathbf{p}^{t}_{N^{t}}]$ contains $N^{t}$ radar points. Each radar point $\mathbf{p}_{i}^{t}$ is characterised by its 3D position $\mathbf{x}_{i}^{t}\in\mathbb{R}^{3}$ in the metric space and auxiliary velocity features $\mathbf{v}_{i}^{t}=[v_{r,i}^{t},v_{c,i}^{t}]$, where $v_{r,i}^{t}$ and $v_{c,i}^{t}$ are the measured relative radial velocity (RRV) and its ego-motion (assumed known) compensated variant. Given each radar point cloud $\mathbf{P^{t}}$ from the stream, our objective is to detect multiple moving objects $\mathbf{D}^{t}=\{\mathbf{d}_{k}^{t}\}_{k=1}^{K^{t}}$ in a class-agnostic manner without the need to regress their 3D bounding boxes. These detected objects are then associated with objects tracked in the previous frame $\mathbf{O}^{t-1}=\{\mathbf{o}_{m}^{t-1}\}_{m=1}^{M^{t}}$. The result of this process is a set of updated objects $\mathbf{O}^{t}$ in track for the current frame.
@@ IV-A Overview
We introduce RaTrack, a generic learning-based framework bespoken for 4D radar-based moving object detection and tracking. As seen in Fig. 1, in our network pipeline, we first apply a backbone (c.f. Sec. IV-B) to encode intra- and inter-frame radar point cloud features. With the extracted features, our point-wise motion estimation module (c.f. Sec. IV-C) infers point-level scene flow as an explicit complement to augment the latent features of radar point clouds. Our advocated idea of class-agnostic detection without bounding boxes is introduced in the object detection module (c.f. Sec. IV-D), in which moving points are first identified and then used to detect moving objects via clustering. Finally, our data association module (c.f. Sec. IV-E) computes the affinity matrix with a learnable distance metric and then optimises the bipartite matching problem. The entire network is end-to-end trainable with a multi-task loss that incorporates three supervised subtasks: motion segmentation, scene flow estimation, and affinity matrix computation.
@@ IV-C Motion Estimation Module
Radar sensors’ inherent sparsity and noise result in latent features from radar point clouds that are deficient in informative geometric cues, complicating object detection and tracking. To address this, we introduce a point-wise motion estimation module that explicitly determines per-point motion vectors. Contrary to the conventional scene flow approach, which estimates per-point motion vectors in a forward direction (from frame $t$ to $t+1$), we opt for a backward estimation (from frame $t$ to $t-1$). This not only negates tracking latency but also ensures that the estimated flow vectors correspond to the points in the current frame $t$.
@@ V-A Evaluation Settings
Baselines. As there are no prior works for 4D radar-based moving object tracking, we select two state-of-the-art LiDAR-oriented 3D MOT methods, i.e., AB3DMOT [10] and CenterPoint [11] as our baselines. To ensure the comparison is fair, we keep their original settings and also train their models on the VoD training split. Note that baseline methods, though designed for LiDAR, also take 4D radar point clouds as input in this work. Specifically, we develop two augmented baselines AB3DMOT-PP and CenterPoint-PP by replacing the detector with PointPillars [52] for AB3DMOT and the backbone with that of PointPillars for CenterPoint.